# Adaptive Query Prompting for Multi-Domain Landmark Detection

### v1

@@ IV-D Ablation Study
We conduct the ablation study to demonstrate the effectiveness of our proposed AQP. Moreover, we compared the results using different vision transformer backbones on the head dataset. Here, we employed three different backbones for testing, namely ViTPose-S, ViTPose-B and ViTPose-L. From the experimental results in Table II, it can be observed that after training with AQP, the model exhibits significant improvements across multiple metrics. Introducing additional prompts effectively increases the trainable parameters, thus enhancing the model’s representational capacity.
TABLE II: Ablation study of our universal model with prompt tuning method using different backbones. ##\## denotes the models are fine-tuned with AQP.
Model MRE($\downarrow$) mm SDR($\uparrow$)(%)

### v2

@@ IV-D Ablation Study
To evaluate the effectiveness of the proposed Adaptive Query Prompting (AQP) method, we conduct ablation experiments on the head dataset using different ViTPose backbones. We compare the performance of models fine-tuned with and without AQP, denoted by “#” in Table II.
As shown in Table II, incorporating AQP consistently improves performance across all backbones and evaluation metrics. This highlights the effectiveness of AQP in adapting a pre-trained model for multi-domain landmark detection. Specifically:
  • Improved MRE: AQP consistently reduces the MRE across all backbones, indicating more accurate landmark localization.
  • Enhanced SDR: AQP significantly improves the SDR, especially at stricter thresholds (2mm, 2.5mm), demonstrating its ability to achieve higher detection accuracy.
These results demonstrate that AQP effectively leverages the pre-trained knowledge from different backbones and adapts it to the specific task of multi-domain landmark detection, leading to improved performance compared to models without AQP.
TABLE II: Ablation study of our universal model with prompt tuning method using different backbones. #\# denotes the models are fine-tuned with AQP.
Model
MRE($\downarrow$)
$m$
SDR($\uparrow$)(%)